OpenStudy (fluttershyk):

The back-to-back stem-and-leaf plot below shows the ages of patients seen by two doctors in a family clinic in one day. Compare the ages of the patients of doctor 1 and doctor 2 using the mean and the median of each data set. Doctor 1 Doctor 2 | 3 | 5 9 2 0 | 2 | 0 0 2 3 6 8 7 5 | 1 | 3 7 9 7 3 2 1 1| 0 | 2 2 3 9 Key: means 29 9|2|3 means 23

OpenStudy (fluttershyk):

@ganeshie8

OpenStudy (rachel.frank40):

For this it would be easiest to start with writing out the ages in a list so that you could then find the mean and median. So you will use the key to tell how to write the ages. For example for Doctor 1 he had seen people of the following ages: 1, 1, 2, 3, 7, 9, 15, 17, 18, 20, 22, 29 Now use the plot to find the ages for Doctor 2. Look at the key.

OpenStudy (fluttershyk):

Okay so Doctor 2 is 35, 20, 20, 22, 23, 26, 13, 17, 2, 2, 3, 9

OpenStudy (rachel.frank40):

Yes, now can you find the mean and median of both of the sets of data?

OpenStudy (fluttershyk):

Yeah I think so

OpenStudy (rachel.frank40):

Remember that mean is where you add up all the numbers and then divide by how many numbers you have. Median is the middle number, if there are an even number of numbers then to find the middle number you will add together those two middle numbers and divide by 2.

OpenStudy (fluttershyk):

Okay thank you

OpenStudy (fluttershyk):

okay @rachel.frank40 heres what I got 1, 1, 2, 3, 7, 9, 15, 17, 18, 20, 22, 29 Doctor 1 35, 20, 20, 22, 23, 26, 13, 17, 2, 2, 3, 9 Doctor 2 9, 15 9+15=24 24/2=12 12 is the median for Doctor 1 1+1+2+3+7+9+15+17+18+20+22+29=144 144/2=72 72 is the mean of Doctor 1 26, 13 26+13=39 39/2=19.5 19.5 is the median for Doctor 2 35+20+20+22+23+26+13+17+2+2+3+9=192 192/2=96 96 is the mean for Doctor 2

OpenStudy (rachel.frank40):

Remember that the mean is when you add and divide by the number of numbers. So for Doctor 1, you have added them up and got 144, but instead of dividing by 2 you are going to divide by 12, because there are 12 people, so the mean for Doctor 1 should be 12. You will also need to change how many you divide by for Dr. 2. Your Median for Dr 1 is correct. But be careful because your list for Dr is not in order from smallest to largest so you need to look at the middle terms.